OpenStudy (anonymous):

Yes, I have this problem to solve. I am confused on how to proceed: 2x-4= 2x+5 Now do I undo by adding a postive 4 from both sides like this? 2x -4 +4 = 2x + 5 + 4 which equals 2x = 2x + 9 after this I am confused... Please help... am I on th right track with this?

OpenStudy (helder_edwin):

yes.

OpenStudy (helder_edwin):

what next?

OpenStudy (helder_edwin):

what would your next step be?

OpenStudy (anonymous):

I subtact 9 from each side to get 2x - 9 + 2x? that is where I am confused. Or do I divide by 2 on each side to get x + x + 9?

OpenStudy (anonymous):

oops x - 9 = x

OpenStudy (helder_edwin):

no. u gotta have the x's on one side.

OpenStudy (anonymous):

did you see 19?

OpenStudy (helder_edwin):

\[ \large 2x=2x+9 \] adding -2x to both sides u get \[ \large 2x-2x=2x-2x+9 \] \[ \large 0=9 \]

OpenStudy (anonymous):

you might start by thinking before you begin is it possible that two times sum number, minus 4, could be the same as two times that same number, plus 5?

OpenStudy (anonymous):

for example, if \(x\) was 10 (which it is not) is it possible that \(2\times 10-4=2\times 10+5\)? or could that work for any number at all? the answer should be more or less clear that if you double the same number, you get the same number, and then if you subtract 4 it cannot be the same as adding 5

OpenStudy (anonymous):

yes that makes sense to me... I was confused becuase I was not sure if I was suppose to, or should I saw allowed to cancel out the varaible...if that makes sense?

OpenStudy (anonymous):

was*

OpenStudy (anonymous):

you can do whatever you want to both sides when you end up with something like \(0=9\) you should realize that there is a problem here, i.e. that you cannot solve this

OpenStudy (anonymous):

thanks a lot... I appreciate your help....I thought that I was totally off on this. I just do not trust my skills! Have a good evening!